OpenStudy (anonymous):

how do i express the set C={x/x+6=10} in roster form

OpenStudy (fedorable):

C = { 4 }

OpenStudy (anonymous):

I thought the elements would contain more counting numbers?

OpenStudy (fedorable):

thats just the answer i got. You could try to investigate my answer more from your own experience with this.

OpenStudy (freckles):

@LauraSmith do you know for what x values x+6 is 10?

OpenStudy (anonymous):

Cool...thanks so much for your help - I was really confused with this one. Sorry to bug, but how did you arrive at that answer?

OpenStudy (freckles):

in other words how do you solve x+6=10

OpenStudy (fedorable):

what is the solution to x+6 = 10

OpenStudy (anonymous):

@ Freckles...Im not really sure sorry. I was hoping it could be explained somehow?

OpenStudy (anonymous):

4

OpenStudy (fedorable):

correct

OpenStudy (fedorable):

that means that C = {x | x+6 = 10} is really the same as C = { 4 }

OpenStudy (anonymous):

That's what I was assuming the answer was, but am not sure how to write that in roster form

OpenStudy (anonymous):

:) awesome

OpenStudy (fedorable):

its already in roaster form, there is only one element in this set because there is only one solution to x+6 = 10

OpenStudy (anonymous):

aha...that's what I was looking for...that explains it all...Thanks so much :)

OpenStudy (fedorable):

np :D

OpenStudy (anonymous):

Fedorable...

OpenStudy (anonymous):

Would you be able to help out with just one more problem - last one

OpenStudy (anonymous):

C ={ X XEN and 7 ≤ X ≤ 14 }

OpenStudy (usukidoll):

is that the set C when \[x \in N \] and 7 ≤ X ≤ 14 ?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

so x is in the set of natural numbers. That inequality is from 7 to 14, so I'm thinking that all the elements from 7 to 14 which is in the interval count as long as it's a natural number. like 8 is a natural number and it's between 7 and 14

OpenStudy (anonymous):

Great, so in other words...it should be C= {7,8,9,10,11,12,13,14}

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

Great, thanks! that's why I was confused with my previous question

OpenStudy (anonymous):

I thought it was more to it...thats all

OpenStudy (fedorable):

oh, sorry i couldn't get to you lol

OpenStudy (anonymous):

No...you were very helpful and I think I was delayed with understanding, sorry

OpenStudy (usukidoll):

what the question is asking is just the set of natural numbers that's between 7 and 14 x in N means x is in natural numbers... natural numbers like 1 2 3 4 5 6.... <--- but we can't write that because it's all less than 7 though watch out some books have 0 counted as a natural number.

OpenStudy (anonymous):

Yes, I was told to ignore 0

OpenStudy (anonymous):

and also fractions

OpenStudy (usukidoll):

I meant the elements in the set of natural numbers that's between 7 and 14. there. This appears in discrete math and some other math subjects yeah 0 doesn't count and for natural numbers it's positive whole numbers.

OpenStudy (anonymous):

hence they are not counting or natural

OpenStudy (usukidoll):

the set of integers has negatives

OpenStudy (anonymous):

agreed!

OpenStudy (anonymous):

speaking of which, I just solved for one :)

OpenStudy (usukidoll):

what math class are you in? just wondering. I've seen this in Discrete Math last semester

OpenStudy (anonymous):

I have Finite Mathematics...it's pretty easy but confusing at times

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

Finite Mathematics and Linear Modeling

OpenStudy (usukidoll):

hmm interesting...

OpenStudy (anonymous):

UsukiDoll

OpenStudy (anonymous):

Do you agree with the answer received for the first problem?

OpenStudy (usukidoll):

yes

OpenStudy (mathmale):

Please post (or bring up) only one question at a time, please. This might bring you faster responses from those who might be able to help you.